# A Survey on Collaborating Small and Large Language Models for Performance, Cost-effectiveness, Cloud-edge Privacy, and Trustworthiness

### v1

@@ Abstract
Large language models (LLMs) have advanced many domains and applications but face high fine-tuning costs, inference latency, limited edge deployability, and reliability concerns. Small language models (SLMs), compact, efficient, and adaptable, offer complementary remedies. Recent work explores collaborative frameworks that fuse SLMs’ specialization and efficiency with LLMs’ generalization and reasoning to meet diverse objectives across tasks and deployment scenarios. Motivated by these developments, this paper presents a systematic survey of SLM–LLM collaboration organized by collaboration objectives. We propose a taxonomy with four goals: performance enhancement, cost-effectiveness, cloud–edge privacy, and trustworthiness. Within this framework, we review representative methods, summarize design paradigms, and outline open challenges and future directions toward efficient, secure, and scalable SLM–LLM collaboration.
@@ 1 Introduction
Large language models (LLMs) have profoundly transformed multiple domains, including AI for science Luo et al. (2022); Al-Lawati et al. (2025); Wang et al. (2024a), programming Shi et al. (2024), and human-centered interaction Zhang et al. (2024c), primarily owing to their massive parameter scales. However, such scale also introduces several challenges: (1) fine-tuning is computationally intensive, limiting efficient model adaptation Thawakar et al. (2025); Liu et al. (2024b); (2) large model size leads to inference latency, constraining real-time applications Leviathan et al. (2023); Kwon et al. (2023); (3) typical edge devices, such as mobile phones, personal computers, and small servers, lack the capacity to host these models, while cloud-based inference raises privacy and cost concerns Carlini et al. (2021); Xu et al. (2024b); and (4) LLMs exhibit inherent reliability risks, including hallucinations and jailbreak vulnerabilities Yao et al. (2024); Farquhar et al. (2024). These challenges underscore the increasing need for customizable, cost-efficient, edge-deployable, and trustworthy AI solutions. Small language models (SLMs), characterized by their compact architecture, low computational cost, and adaptability, have emerged as promising counterparts to mitigate these issues, although their general reasoning and knowledge coverage remain more limited than those of LLMs. Consequently, harnessing the complementary strengths of SLMs and LLMs provides a viable pathway toward developing efficient, scalable, and reliable AI systems.
@@
Researchers have proposed diverse approaches for SLM–LLM collaboration, leveraging SLMs’ strengths in customization, efficiency, and local deployment alongside LLMs’ generalization and reasoning capabilities Xu et al. (2024a); Chen et al. (2024); Wang et al. (2025b). Despite notable progress, a systematic survey focused on the objectives of SLM–LLM collaboration remains absent. Existing studies largely pursue four objectives. (1) Performance: integrating domain-specific SLMs with general LLMs to improve performance across specialized and general tasks. (2) Cost-effectiveness: using SLMs for lightweight processing and invoking LLMs selectively to reduce computational and API costs. (3) Cloud–edge privacy: employing on-device SLMs for private data handling while cloud LLMs offer broader reasoning to balance efficiency and privacy. (4) Trustworthiness: adopting SLMs as safety policy encoders that guide LLMs to produce safer and more reliable outputs. This paper presents the first comprehensive survey around the objectives of SLM–LLM collaboration: performance, cost-effectiveness, cloud–edge privacy, and trustworthiness. We propose a taxonomy, summarize representative works (Tab.1 in Appendix), and outline future directions.
@@
Several surveys have examined SLMs in the LLM era. Wang et al. (2024b, 2025a) provided broad overviews of SLM design, applications, and reliability but mentioned collaboration only briefly. Lu et al. (2024); Van Nguyen et al. (2024); Xu et al. (2024b) analyzed SLM advantages, architectures, and deployment, yet offered limited discussion of collaboration. Existing collaboration surveys, such as Chen and Varoquaux (2024); Niu et al. (2025); Li et al. (2025b), focus on one-way or cloud–edge settings without addressing broader objectives. To fill this gap, we systematically review SLM–LLM collaboration by collaboration objectives, offering key insights to guide future studies.
@@ 2 Collaboration for Performance
The optimal LLM varies across tasks and queries Jiang et al. (2023), making single-model solutions suboptimal. Leveraging complementary models is therefore effective. Owing to their customizability, SLMs have produced many domain-specific products, such as SciGLM Zhang et al. (2024a), ChemLLM Zhang et al. (2024b), and Biomistral Labrak et al. (2024). Hence, SLM–LLM collaboration for performance has been a prevalent strategy. We divide existing works mainly into two paradigms: (1) Guidance–Generation, where one model steers another’s generation, and (2) Division–Fusion, where the SLMs and LLMs leverage their complementary strengths on distinct tasks.
@@ 2.1 Guidance–Generation Collaboration
As shown in Fig. 1(a), the LLM uses its semantic understanding to clarify tasks and provide fine-grained guidance for task-specific SLMs. For example, SynCID Liang et al. (2024) employs LLM-generated task descriptions to guide SLM reasoning.
@@
As shown in Fig. 1(b), the SLM offers domain expertise and cues, while the LLM integrates this information for more accurate and reliable outputs. For instance, SuperICL and its variant Xu et al. (2024a); Mojarradi et al. (2024) inject SLM predictions and confidence scores into the LLM’s context, and LM-Guided CoT Lee et al. (2024) uses SLM-generated reasoning chains to guide LLM inference.
@@ 2.2 Division–Fusion Collaboration
Division–fusion collaboration paradigm is preferred when multiple models exhibit heterogeneous capabilities and no single backbone model is suitable. According to different task needs, this paradigm mainly includes (1) Parallel Ensemble for collective intelligence and (2) Sequential Cooperation for multi-stage workflows.
@@
As shown in Fig. 2, multiple SLMs and LLMs work in parallel, and their outputs are integrated for higher accuracy. Two common strategies are majority voting, as in ELM Gondara et al. (2025), which aggregates outputs by consensus, and cross-verification, as in CaLM Hsu et al. (2024), which iteratively refines results until consensus is reached.
Figure 2: Parallel Ensemble: Collective Intelligence
@@
For multi-stage tasks, subtasks are assigned to suitable models and connected in a pipeline, as shown in Fig. 3. When tasks are explicitly staged, SLMs and LLMs take different roles: SLMs handle precise components (e.g., schema matching in ZeroNL2SQL Fan et al. (2024) and KDSL Chen et al. (2025a)), while LLMs manage complex reasoning (e.g., GCIE Bao and Wang (2024)). When stages are implicit, the LLM acts as a planner and SLMs as executors, as in HuggingGPT Shen et al. (2023) and TrajAgent Du et al. (2024), which assign subtasks to expert SLMs and integrate them.
Figure 3: Sequential Cooperation: Division of Labor

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable progress across domains and applications, but face challenges such as high fine-tuning costs, inference latency, limited edge deployability, and reliability concerns. Small language models (SLMs), with compact, efficient, and adaptable features, offer promising solutions. Building on this potential, recent research explores collaborative frameworks that integrate their complementary strengths, leveraging SLMs’ specialization and efficiency with LLMs’ generalization and reasoning, to address diverse objectives across tasks and deployment scenarios. Motivated by these developments, this paper presents a systematic survey of SLM–LLM collaboration from the perspective of collaboration objectives. We propose a taxonomy covering four goals: performance enhancement, cost-effectiveness, cloud–edge privacy, and trustworthiness. Under this framework, we review representative methods, summarize design paradigms, and outline open challenges and future directions toward efficient and secure SLM-LLM collaboration. The collected papers are available in link here.
@@ 1 Introduction
Figure 1: Detailed taxonomy of SLM–LLM collaboration, depicting the sub-taxonomy under the collaboration objectives of performance, cost-effectiveness, privacy, and trustworthiness.
@@
Large language models (LLMs) have profoundly transformed multiple domains, including AI for science Luo et al. (2022); Al-Lawati et al. (2025); Wang et al. (2024a, d), programming Shi et al. (2024), and human-centered interaction Zhang et al. (2024c), primarily owing to their massive parameter scales. However, such scale also introduces several challenges: (1) fine-tuning is computationally intensive, limiting efficient model adaptation Thawakar et al. (2025); Liu et al. (2024d); Xia et al. (2024); Zhang et al. (2024h); (2) large model size leads to inference latency, constraining real-time applications Leviathan et al. (2023); Kwon et al. (2023); Cai et al. (2024); (3) typical edge devices, such as mobile phones, personal computers, and small servers, lack the capacity to host these models, while cloud-based inference raises privacy and cost concerns Carlini et al. (2021); Xu et al. (2024c); Li et al. (2024b); and (4) LLMs exhibit inherent reliability risks, including hallucinations and jailbreak vulnerabilities Yao et al. (2024); Farquhar et al. (2024); Wang et al. (2023a). These challenges underscore the increasing need for customizable, cost-efficient, edge-deployable, and trustworthy AI solutions. Small language models (SLMs), characterized by their compact architecture, low computational cost, and adaptability, have emerged as promising counterparts to mitigate these issues, although their general reasoning and knowledge coverage remain more limited than those of LLMs. Consequently, harnessing the complementary strengths of SLMs and LLMs provides a viable pathway toward developing efficient, scalable, and reliable intelligent systems.
SLM-LLM Collaboration Performance §2 Guidance-Generation §2.1 SynCIDLiang et al. (2024), PiFiKim et al. (2025), SuperICLXu et al. (2024a), Ensemble-SuperICLMojarradi et al. (2024), LM-Guided CoTLee et al. (2024), CoWestJiao et al. (2024), Synergistic AugmentationLi et al. (2025d), MSLLMQiao et al. (2025), Flipping KD Li et al. (2025b) Division-Fusion §2.2 ELMGondara et al. (2025), CaLMHsu et al. (2024), ZeroNL2SQLFan et al. (2024), KDSLChen et al. (2025a), Collab-RAG Xu et al. (2025), GCIEBao and Wang (2024), HuggingGPTShen et al. (2023), TrajAgentDu et al. (2024) Cost-Effectiveness §3 Pre-Training Stage §3.1 SALTRawat et al. (2024), Perplexity-Based Data PruningAnkner et al. (2024), LiGOWang et al. (2023b), Progressive TrainingYano et al. (2025),HyperCloning Samragh et al. (2024), DistilBERTSanh et al. (2019), TinyBERTJiao et al. (2020), MiniLM Wang et al. (2020), Pre-training DistillationPeng et al. (2025), DPTGoyal et al. (2025), MINIPLMGu et al. (2025) Fine-Tuning Stage §3.2 LLKDLi et al. (2025a), LLM-to-SLM KDJazbec et al. (2024), HAWKEYEShe et al. (2025), Causal DistillationMuhebwa and Osman (2025), C2KDChen et al. (2025d), Agent Distillation Kang et al. (2025), LSC4RecLv et al. (2025), SmallPlanPham et al. (2025), LLM-to-SLMBergner et al. (2024), LoRAMZhang et al. (2025a), LiteMOEZhuang et al. (2024), GateKeeperRabanser et al. (2025), Logit-fusionFan et al. (2025c) Inference Stage §3.3 FrugalGPTChen et al. (2024a), Yue et al. (2024), P3DeferZhang et al. (2024d), RouteLLM Ong et al. (2025), SlimPMTan et al. (2024), Division-of-thoughtsShao et al. (2025), R2RFu et al. (2025b), CITERZheng et al. (2024), MixLLMWang et al. (2025c), MoE2Jin et al. (2025), BeagleZhong et al. (2025), LookAhead ReasoningFu et al. (2025c), APDIsrael et al. (2025), LLMCADXu et al. (2023), Edge–cloud HybridHao et al. (2024), PICEZhan et al. (2025), FS-GENZhang et al. (2024e), BestRouteDing et al. (2025), AgentTTSWang et al. (2025b) Cloud-Edge Privacy §4 Inference Stage §4.1 CasperChong et al. (2024), LLM Gatekeeper Uzor et al. (2025), LLM-AnonymizerWiest et al. (2024), Hartmann et al. (2024), COREFan et al. (2025a), P3DeferZhang et al. (2024d), RemoteRAGCheng et al. (2025), CoGenesisZhang et al. (2024f), PrivacyBoost-SLMZhang et al. (2024i), Apple IntelligenceGunter et al. (2024) Fine-tuning Stage §4.2 MiniLLMGu et al. (2024), LlamaDuoPark et al. (2025), DRAGChen et al. (2025c), HomeLLaMAHuang et al. (2025c), On-device PersonalizationQin et al. (2024), TinyLLMTian et al. (2025), Mix DistillationLi et al. (2025g), ADPAGao et al. (2025b), CombLMOrmazabal et al. (2023), Logit OffsetsLiu et al. (2024a); Wang et al. (2025d), CPTHe et al. (2024), GradOTYao et al. (2025), GOODFang et al. (2024a), PHLoRAVasani et al. (2025), LoRASuiteLi et al. (2025f), LoRA-XFarhadzadeh et al. (2025), Trans-LoRAWang et al. (2024c), Cross-LoRAXia et al. (2025), CrossLMDeng et al. (2025), LSRPZhang et al. (2025c), FedMKTFan et al. (2025b), FedPTGao et al. (2024), FedPFTPeng et al. (2024), FDLoRAQI et al. (2024), FedPipeFang et al. (2024b), FLoRAWang et al. (2024e) Trustworthiness §5 Safety-Guided Decoding §5.1 Purifying LLMsLi et al. (2024a), MODShi et al. (2024), Logit offsetMitchell et al. (2024), Offset UnlearningHuang et al. (2025a), Weak-to-Strong JailbreakZhao et al. (2025) SLM as Guardian §5.2 Llama GuardInan et al. (2023); Team (2024); Dubey et al. (2024); Fedorov et al. (2024); Chi et al. (2024); AI (2025a), Prompt GuardAI (2024, 2025b), ShieldGemmaZeng et al. (2024, 2025), WildGuardHan et al. (2024), ThinkGuardWen et al. (2025), MiniCheckTang et al. (2024), GuardrailsNagireddy et al. (2024), LlamaFirewallChennabasappa et al. (2025)
Figure 2: A Taxonomy of SLM-LLM Collaboration.
@@
Researchers have developed various approaches to enable collaboration between SLMs and LLMs, leveraging the strengths of SLMs in customization, efficiency, and local deployment, alongside the powerful generalization capabilities of LLMs Xu et al. (2024a); Chen et al. (2024a); Wang et al. (2025b); Gu et al. (2024); Inan et al. (2023). Despite the importance of this topic and the various approaches developed, there is a lack of a systematic survey centered on the objectives of SLM–LLM collaboration. Hence, we fill this gap by providing a taxonomy that unifies recent works under clear collaboration objectives. Existing studies generally pursue four main objectives. First, by combining domain-specific SLMs that outperform general-purpose LLMs in specialized contexts, many works target collaboration for enhanced task performance across both general and domain-specific tasks. Second, emphasizing cost-effectiveness, SLMs are employed for lightweight processing, while LLMs are invoked selectively when necessary. Third, under the paradigm of cloud–edge collaboration, on-device SLMs process privacy-sensitive data, whereas cloud-based LLMs provide general reasoning, ensuring both effectiveness and privacy. Finally, SLMs act as lightweight safety policy encoders, while LLMs serve as backbone generators to improve output trustworthiness. Therefore, our survey is organized around the objectives of SLM–LLM collaboration: performance, cost-effectiveness, cloud–edge privacy, and trustworthiness. Under this framework, we propose a taxonomy of existing studies (in Fig. 2 and Tab. 1), summarize key insights, and outline future directions for efficient, secure, and deployable intelligent systems.
@@
Several surveys have examined SLMs in the LLM era. Wang et al. (2024b, 2025a) provided a broad overview of SLM design, acquisition, applications, and trustworthiness but mentioned SLM–LLM collaboration only briefly. Lu et al. (2024) evaluated SLM advantages through experimental analysis, Van Nguyen et al. (2024) focused on architectures and compression techniques, and Xu et al. (2024c) discussed on-device deployment, all with limited discussion of collaboration. Existing collaboration surveys also remain narrow: Chen and Varoquaux (2024) explored one-way collaboration (SLMs for LLMs or LLMs for SLMs), whereas Niu et al. (2025); Li et al. (2025c) examined cloud–edge collaboration without addressing diverse collaboration objectives. To fill this gap, we systematically review existing studies from the perspective of collaboration objectives, summarizing key insights to guide future SLM–LLM collaboration.
Our main contributions are summarized as follows: (1) We present the first systematic survey of SLM–LLM collaboration from a novel perspective centered on collaboration objectives, encompassing performance, cost-effectiveness, cloud–edge privacy, and trustworthiness. The overall taxonomy is illustrated in Fig. 2. (2) We analyze and categorize performance-oriented collaboration studies based on how they exploit the complementary strengths of SLMs and LLMs, forming two major paradigms: Guidance–Generation and Division–Fusion. (3) We summarize cost-effectiveness-driven collaboration along the lifecycle of language models, covering the pre-training, fine-tuning, and inference stages. (4) We review privacy-preserving collaboration between cloud and edge models according to how edge data are privately leveraged by cloud models during both the inference and fine-tuning stages. (5) We discuss trustworthiness-oriented collaboration under two paradigms: safety-guided decoding and guardian–generator frameworks.
I. Collaboration for Performance
LLM guided SLM PiFi (2025) BERT, T5 LLaMA 3.1-8B SST-2, IMDB, etc. Frozen LLM layer enhances SLM
SLM guided LLM CoWest Jiao et al. (2024) LLaMA-3-8B GPT-4/3.5 IfQA, MedMCQA, etc. Preference-aligned weak-to-strong collab.
[…]
Sequential KDSL (2025a) Qwen2-VL 2B Qwen2-VL 7B SLM rules; LLM inference
II. Collaboration for Cost-effectiveness
Data curation Ankner et al. (2024) MPT-125M MPT-1B/3B The Pile, Dolma, etc. SLM perplexity-guided efficient pruning
Co-curriculum LiGO (2023b) BERT-S, DeiT-S BERT-B, DeiT-B GLUE, SQuAD, ImageNet Learned growth reduces training cost
Co-curriculum HyperCloning (2024) OPT-350M OPT-1.3B DOLMA, The Pile, etc. Function-preserving SLM-to-LLM initial.
Pre-train distill. DPT (2025) Synthetic bigram data Distillation improves scaling, harms ICL
Pre-train distill. MiniPLM (2025) Qwen 200M/500M Qwen1.5-1.8B DCLM subset Offline sampling accelerates pretraining
Selective KD LLM-to-SLM (2024) T5-Small T5-Large WMT, CNN, etc. Frozen LLM conditions fast SLM
Proxy transfer GateKeeper (2025) Gemma-2B Gemma-7B CIFAR, ARC, VQAv2, etc. Conf-tuned SLM improves cascading
Proxy transfer G-Boost (2025c) TinyLlama-1B LLaMA2-13B GSM8K, MATH-500 Reward-guided dynamic collaboration
Cascade Routing R2R (2025b) DeepSeek-R1-1.5B DeepSeek-R1-32B AIME, GPQA Token-level routing for efficiency
Cascade Routing SlimPM (2024) LLaMA2-7B LLaMA2-70B NQ, TriviaQA, etc. Proxy-guided selective retrieval efficiency
Cascade Routing CITER (2024) Qwen2-1.5B Qwen2-72B CSQA, ARC, etc. Token-level routing boosts efficiency
Cascade Routing MixLLM (2025c) Llama-3-8B GPT-3.5/4, etc. RouterBench Adaptive routing
Speculative APD (2025) Qwen2.5-0.5B Dream-7B GSM8K, GPQA, MATH Adaptive parallel decoding via PoE
Speculative LLMCAD (2023) GPT-2/mT5 Vicuna/LLaMA2 CNN/DailyMail, etc. Token-tree drafting
Speculative Cloud-edge (2024) TinyLlama GPT-3.5 GSM8K Token-level edge–cloud collaboration
Optimal TTS AgentTTS (2025b) LLaMA-3-3B LLaMA-3-70B HotpotQA, CWQ, etc. Budget-aware collaboration
Optimal TTS BestRoute (2025) Phi-3-mini/medium, etc. GPT-4o/3.5 MixInstruct, RewardBench, etc. Adaptive routing minimizes cost
III. Privacy-aware Cloud–Edge Collaboration
PII Gatekeeper LLM-Anonymizer (2024) Llama-3 8B, etc. Any real clinical letters Local LLM anonymization
PII Gatekeeper P3Defer (2024d) Gemma-2B Gemma-7B GSM8K, MedSum, etc. CoT policy RL reduces leakage
All-info Guardian PrivacyBoost-SLM (2024i) BioLinkBERT, etc. GPT-3.5 MedQA, HEADQA, etc. Privacy prompting boosts medical QA
SLM as Learner MiniLLM (2024) GPT-2-125M GPT-2-1.5B DollyEval, SelfInst, etc. Distill LLM into SLM
SLM as Learner Mix Distillation (2025g) Qwen2.5-0.5B, etc. QwQ-32B, etc. MATH, GSM8K, etc. Learnability gap, mixed distillation
LLM as Learner Proxy-tuning (2024a) LLaMA-2-7B LLaMA-2-70B GSM, Toxigen, etc. Steer LLM with private SLM
LLM as Learner Cross-LoRA (2025) Qwen2.5-1.5B/3B, etc. Hetero-families ARC-c, ARC-e, OBQA, etc. Zero-data subspace LoRA transfer
Co-learners FedMKT (2025b) GPT-2-xl, OPT, etc. LLaMA2-7B RTE, WIC, BoolQ, etc. Bidirectional federated co-distillation
IV. Collaboration for Trustworthiness
Safe decoding LLM Purifying (2024a) CodeLLaMA-7B CodeLLaMA-13B HumanEval, LAMBADA, etc. Logit fusion for clean gen.
@@ 2 Collaboration for Performance
The optimal LLM varies across tasks and queries Jiang et al. (2023); Ashiga et al. (2025); Bell (2025); Wang et al. (2025b), making single-model solutions suboptimal. Leveraging complementary models is therefore effective. Owing to their customizability, SLMs have produced many domain-specific products, such as SciGLM Zhang et al. (2024a), ChemLLM Zhang et al. (2024b), and Biomistral Labrak et al. (2024). Hence, SLM–LLM collaboration for performance has been a prevalent strategy. We divide existing works mainly into two paradigms: (1) Guidance–Generation, where one model steers another’s generation, and (2) Division–Fusion, where the SLMs and LLMs leverage their complementary strengths on distinct tasks.
@@ 2.1 Guidance–Generation Collaboration
As shown in Fig. 3(a), the LLM uses its semantic understanding to provide fine-grained guidance for task-specific SLMs. For example, SynCID Liang et al. (2024) employs LLM-generated task descriptions to guide SLM reasoning. PiFi Kim et al. (2025) uses LLMs’ certain knowledge layer plug-in to enhance SLM’s generalization and understanding.
@@
As shown in Fig. 3(b), the SLM offers domain expertise and cues, while the LLM integrates this information for more accurate and reliable outputs. For instance, SuperICL and its variant Xu et al. (2024a); Mojarradi et al. (2024) inject SLM predictions and confidence scores into the LLM’s context, and LM-Guided CoT Lee et al. (2024) uses SLM-generated reasoning chains to guide LLM inference. CoWest Jiao et al. (2024) aligns weak–strong preferences to synergize domain expert SLMs’ knowledge and LLMs’ general reasoning. Synergistic Augmentation Li et al. (2025d) uses SLM to guide LLM with confident demonstrations and slot hints, boosting zero-shot slot filling performance. MSLLM Qiao et al. (2025) combines SLM’s precise error detection with LLM’s fluency and semantic understanding for superior Chinese spelling correction. Flipping KD Li et al. (2025b) enables LLMs to learn task-specific ability from SLMs.
@@ 2.2 Division–Fusion Collaboration
Division–fusion collaboration paradigm is preferred when multiple models exhibit heterogeneous capabilities and no single backbone model is highly suitable. According to different task needs, this paradigm mainly includes (1) Parallel Ensemble for collective intelligence and (2) Sequential Cooperation for multi-stage workflows.
@@
As shown in Fig. 4, multiple SLMs and LLMs work in parallel, and their outputs are integrated for higher accuracy. Two common strategies are majority voting and cross-verification. For majority voting, as in ELM Gondara et al. (2025), multiple independent solutions are sampled from a set of distinct models and combined by consensus to integrate multi-model capabilities. For cross-verification, as in CaLM Hsu et al. (2024), an LLM and an SLM independently generate answers in parallel and iteratively refine both inputs and model states, continuing the process until their outputs reach consensus.
Figure 4: Parallel Ensemble: Collective Intelligence.
@@
For multi-stage tasks, subtasks are assigned to suitable models and connected in a pipeline, as shown in Fig. 5. When tasks are explicitly staged, SLMs and LLMs take different roles: SLMs handle precise components (e.g., schema matching in ZeroNL2SQL Fan et al. (2024) and KDSL Chen et al. (2025a), question decomposition in Collab-RAG Xu et al. (2025)), while LLMs manage complex reasoning (e.g., GCIE Bao and Wang (2024)). When stages are implicit, the LLM acts as a planner and SLMs as executors, as in HuggingGPT Shen et al. (2023) and TrajAgent Du et al. (2024), which assign subtasks to expert SLMs and integrate them.
Figure 5: Sequential Cooperation: Division of Labor.